O 28-17: Amortized Loans Problem 28-30 Loan Amortization Your company is planning to borrow $2.25 million on a 5-year, 12%, annual payment, fully amortized term loan, what fraction of the payment made at the end of the second year will represent repayment of principal? Round your answer to two decimal places. 1%
Answer #1

To calculate Principal , interest part of the Periodic Payment we need to create amortization table

Beginning Balance= 2,250,000 PMT =PMT(12%,5,-2,250,000) Interest part of PMT=12%*Beginning Principal Principal part of PMT= PMT-Interest Ending Balance= Beginning balance-Principal part of PMT
1 ₹ 2,25,000.00 ₹ 62,417.19 ₹ 27,000.00 ₹ 35,417.19 ₹ 1,89,582.81
2 ₹ 1,89,582.81 ₹ 62,417.19 ₹ 22,749.94 ₹ 39,667.25 ₹ 1,49,915.56
3 ₹ 1,49,915.56 ₹ 62,417.19 ₹ 17,989.87 ₹ 44,427.32 ₹ 1,05,488.24
4 ₹ 1,05,488.24 ₹ 62,417.19 ₹ 12,658.59 ₹ 49,758.60 ₹ 55,729.63
5 ₹ 55,729.63 ₹ 62,417.19 ₹ 6,687.56 ₹ 55,729.63 ₹ 0.00

Fraction of payment in second year represented by principal = 39667.25/62417.19 = 0.6355 or 63.55%


Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well